PartnerObserving Log for 2007 April 17/18

Observers: J. Harris, D. Sand
Instrument Support: D. Thompson, R. Bertram
Telescope Operator:
Instrument: LBC-Blue

$ 11:00: Early in the day, we tested the camera shutter with different filters and rotation angles; it seems to be working. However, we have only tested it with the telescope pointing at the horizon so far.

$ 16:30: Taking biases and 900-sec dark.

The dome shutter and the camera shutter are both operational at the moment. We're going to do more tests of the camera shutter before sunset, including tests with the telescope pointed at zenith.

At zenith, after 5 exposures we got an RPC communication error, requiring a CMU reboot. We don't think this is related to the shutter, it just happens sometimes.

After we resolved this issue, we started getting staurated exposures, leading us to suspect the shutter was stuck open. We confirmed this by taking a Dark, which was also saturated.

Next, we returned the telescope to the horizon (where last night's tests showed no shutter issues), but even at this position, we could not get the shutter un-stuck. Ray is now going to exercise the shutter by going directly up to the camera.

Hours passed. Dave and Ray took LBC off the telesscope and into the lab. John Hill arrives from Tucson at 1ish.

John and Ray disassembled the shutter and fiddled with things, but never really found anything wrong. Shutter works OK in the lab unless you stall it with your finger. We tested in the lab in both horizon-like and zenith-like orientations. Put cryostat back on the telescope at ~3:30.

Took several biases and test images pointed at the horizon. Now moving to zenith at 3:42.

$ 4:15: LBC UI hung after stopping an OB while telescope wasn't running. lbckill/lbcstart

$ 4:24: After focusing, started sdt_fan1_U_6

First image was egg-shaped. Ugh. Apparently this was coma that LBCFPIA didn't correct in the first two iterations. Seeing is varying between 1 and 2 arcsec.

$ 4:30: Abort during the middle of position 2 of the dither sequence. Trying to refocus. LBCFPIA added 500 nm of Z8.

Trying to start sdt_fan1_U_6 again at 4:38. Mostly just to see if focus is better. We only took one image. Focus is better but seeing still variable up to 2 arcsec. Going to do sky flats now.

$ 4:55: Taking B sky flats (t=30 sec). Counts 7500--20,000

$ 5:03: Taking Uspec sky flats (t=30 sec). Counts 10,000--23,000

$ 5:14: Taking U-Bessel flats (t=10 sec). Counts 17,000--40,000

$ 05:18: Taking more U-Bessel flats (t=3 sec). 18,000--37,000

Closed Dome.

$ 05:38: Move telescope back to zenith for more shutter testing. Taking a series of 30 sec V expoures.

-- JasonHarris - 17 Apr 2007
Topic revision: r6 - 21 Jul 2009, NormCushing
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback